For HA: µ < 45, s=5, n=35, at ? = 0.01, what is the critical value of the test statistic with 3 decimal places?
Answer #1

Given that, HA: µ < 45, s=5, n=35

and significance level (\alpha) = 0.01

Degrees of freedom (DF) = n - 1 = 35 - 1 = 34

Since it is one-tailed ( left-tailed ) test, t-critical value at significance level (\alpha) = 0.01 with DF = 34 is tcrit = -2.441

Hence, critical valur of the test statistic is -2.441

Note: since population standard deviation is unknown we used t-table for finding critical value.